OpenStudy (anonymous):

The point (12, 5) is on the terminal side of an angleθ. What is cos θ? a. 12 b. 12/5 c. 12/13 d. 13/12

OpenStudy (kc_kennylau):



OpenStudy (kc_kennylau):

Do you remember the definition for cosine?

OpenStudy (anonymous):

cos(a)= b/c

OpenStudy (anonymous):

right?

OpenStudy (kc_kennylau):

Yep, and in this case?

OpenStudy (anonymous):

cos(12)=5/13??

OpenStudy (kc_kennylau):

You don't apply cosine to a side; you apply it to an angle

OpenStudy (kc_kennylau):

That means cos(θ) instead of cos(12)

OpenStudy (kc_kennylau):

And you applied it to the wrong sides; it should be 12/13

OpenStudy (anonymous):

so cos(θ)=12/13

OpenStudy (kc_kennylau):

Yep